The graph plotted between KE of a particle with its momentum is:
A. Straight line
B. Hyperbola 
C. Parabola
D. Circle

Grade:12th pass

1 Answers

Arun
4 years ago

E=1/2mv^2

2E.m=(m.v)^2

Since, P=m.v

2.E.m=p^2

√E =p/(√2m)

 

Hence Graph is parabolic

Option C is correct